OpenStudy (anonymous):

Create your own unique quadratic equation in the form y = ax2 + bx + c. Use complete sentences and show all work to determine the following: Does the graph open up or down? How do you know? Explain whether the graph has a maximum or minimum point. Find the vertex and x-intercepts of the graph.

OpenStudy (anonymous):

MEDALS FOR ALL pleaseeeeee

OpenStudy (anonymous):

are you still there ?

OpenStudy (bahrom7893):

okay, let's make up an easier equation lol to make our lives easier: Let a = 2, b = 4, c = 4

OpenStudy (anonymous):

lol okay... if you can do this problem for me, i promise not to post another... i just really have to bring my grade up.... 58 F right now :(

OpenStudy (bahrom7893):

wait let a = 4

OpenStudy (bahrom7893):

Hey i didn't say u werent allowed to post anymore.

OpenStudy (anonymous):

i know :( its just i feel like im annoying yall... i just really need to bring that F up...

OpenStudy (bahrom7893):

so, here's our equation a=b=c=4 y = 4x^2+4x+4, very easy to factor lol

OpenStudy (bahrom7893):

It opens up because a is positive (4>0)

OpenStudy (bahrom7893):

Graph has a minimum point at it's turning point which is when x=-b/(2a) = -4/8 = -1/2

OpenStudy (bahrom7893):

So min: (-1/2;9/4)

OpenStudy (bahrom7893):

vertex is the same as minimum so (-1/2; 9/4). x intercept occurs when y=0: 0 = (2x+2)^2, so 2x+2 = 0; 2x=-2; x=-1 (-1;0) is the x intercept.

OpenStudy (bahrom7893):

it has only one x intercept, there done lol

OpenStudy (anonymous):

I really appreciate this.. your help probably just boosted my grade. I could've never done all that... I really do appreciate it. you made my day. THANK YOUUU :)))

OpenStudy (bahrom7893):

awwww, np :)

OpenStudy (anonymous):

& My grade flew up :))))

OpenStudy (bahrom7893):

awesome!

OpenStudy (anonymous):

You're very intelligent!

OpenStudy (bahrom7893):

nah, i just learned this a long time ago lol

OpenStudy (anonymous):

Oh.... lol